# Searching for the tetraneutron resonance on the lattice

### v1

@@ Abstract
The nature of the tetraneutron ($4n$) system remains a pivotal question in nuclear physics. We investigate the $4n$ system using nuclear lattice effective field theory in finite volumes with a lattice size up to $L=30$ fm, employing both a high-precision N3LO interaction and a simplified SU(4) symmetric one. The ground-state energy is found to decrease smoothly with increasing box size, showing no plateau characteristic of a resonance. We further compute the dineutron-dineutron scattering phase shift using Lüscher’s finite-volume method. The phase shift is negative at low momenta, indicating repulsion in the dilute limit. At intermediate momenta, it exhibits a weak attraction with a peak of approximately $10^{\circ}$ at relative momentum of 60-84 MeV. While this structure does not constitute a resonance, the corresponding confined $4n$ energy of 1.7–3.3 MeV lies close to the experimentally observed low-energy peak.
@@ I Introduction
Bound nuclear systems conventionally form when protons and neutrons coalesce through the strong interaction. However, the combination of attractive strong interactions and the absence of Coulomb repulsion makes the existence of pure neutron nuclei a plausible concept, a quest that dates back to the early 1960s Ogloblin and Penionzhkevich (1989). Confirming or ruling out the existence of such neutral nuclei, whether as bound or resonant states, is crucial for advancing our understanding of nuclear physics, the strong interaction, universal fermion properties in the unitary limit Elhatisari et al. (2017); Braaten and Hammer (2006), and astrophysical objects Matsuki et al. (2025); Ivanytskyi et al. (2019).
@@
While a bound $4n$ nucleus is generally believed to be excluded Lazauskas et al. (2023); Marqués and Carbonell (2021); Bertulani and Zelevinsky (2003), the interpretation of these signals as a resonance remains highly debated. A variety of theoretical approaches, including Jost function analysis in the complex momentum plane Sofianos et al. (1997), the hyperspherical harmonic method Grigorenko et al. (2004), Faddeev-Yakubovsky equations Lazauskas and Carbonell (2005); Hiyama et al. (2016); Carbonell et al. (2017), the Gaussian expansion method Hiyama et al. (2016); Carbonell et al. (2017), the no-core Gamow shell model and density matrix renormalization group Fossez et al. (2017), the $4n$ response function Lazauskas et al. (2017), Alt, Grassberger, and Sandhas (AGS) equations Deltuva (2018); Deltuva and Lazauskas (2019b), and adiabatic hyperspherical methods Higgins et al. (2020, 2021), converge on the absence of a $4n$ resonance. The low-energy peak observed by Duer and others (2022) has been explained as a consequence of dineutron-dineutron correlations through a reaction model study Lazauskas et al. (2023). Reactions with multi-neutron final states using a nonrelativistic conformal field theory are investigated in Hammer and Son (2021). Recent tensor-optimized antisymmetrized molecular dynamics studies suggest that a $4n$ resonance at the nuclear surface may be influenced by core attraction Wan et al. (2025). For further insight into multi-neutron correlations in light nuclei, see Zhang et al. (2025b).
@@
In contrast, many-body methods that employ an external potential trap to confine the $4n$ system, subsequently extrapolating to the no-trap limit, tend to support the existence of a $4n$ resonance with energies and widths close to the experimental values. These include quantum Monte Carlo Pieper (2003); Gandolfi et al. (2017), the no-core shell model Shirokov et al. (2016), and the no-core Gamow shell model Li et al. (2019). However, as discussed in Refs. Deltuva and Lazauskas (2019a); Ishikawa (2020), threshold effects are critical and must be carefully accounted for in such extrapolations. For a comprehensive review of previous efforts, see Marqués and Carbonell (2021).
@@
To further elucidate the nature of the $4n$ signal reported in Duer and others (2022), additional theoretical investigations are needed. In this work, we study the $4n$ system using nuclear lattice effective field theory (NLEFT) Lee (2025); Lähde and Meißner (2019), a method that has proven very successful in providing a unified description of nuclear phenomena on a discretized lattice, including scattering Elhatisari et al. (2025); Wu et al. (2025); Elhatisari et al. (2017, 2015), clustering Epelbaum et al. (2011, 2012); Shen et al. (2023, 2025), nuclear matter Tong et al. (2025a, b); Ma et al. (2024); Ren et al. (2024); Lu et al. (2020), and the structure of light and heavy nuclei Hildenbrand et al. (2025); Ren et al. (2025); Song et al. (2025); Niu and Lu (2025); Wang et al. (2025); Zhang et al. (2025a); Hildenbrand et al. (2024); Meißner et al. (2024); Sarkar et al. (2023); Lu et al. (2022); Shen et al. (2021). We calculate the ground-state energy of $4n$ in finite volumes up to a lattice size of $L\simeq 30$ fm. No external trap is applied, as the finite volume itself provides the necessary confinement for the dilute system.
@@ II Formalism
with $\mathbf{n}=(n_{x},n_{y},n_{z})$ is the lattice coordinate vector, with each component ranging from 0 to L1L-1, and $\mathbf{k}$ is the lattice momentum vector. The initial state $|\Psi_{0}\rangle$ is constructed as a Slater determinant from the two lowest momentum states, each with spin up and down. Furthermore, we perform an angular momentum projection onto the $A_{1}^{+}$ irreducible representation of the cubic group to project out the $0^{+}$ state and accelerate the Euclidean time evolution Johnson (1982); Lu et al. (2014). The ground state energy at a given projection time $\tau$ is calculated as
@@
It is well known that the tetraneutron system lacks bound subsystems. However, when confined in a finite volume, the dineutron becomes quasi-bound. Following the method for dimer-dimer scattering on a lattice Elhatisari et al. (2017), we compute the 2n2n-$2n$ scattering phase shifts using Lüscher’s formula. It relates the two-body energy levels in a periodic cubic box to the elastic scattering phase shifts Luscher (1986, 1991):
@@
with $S(\eta)$ the regulated three-dimensional zeta function
$S(\eta)=\lim_{\Lambda\to\infty}\left[\sum_{\vec{n},\,|\vec{n}|\leq\Lambda}\frac{1}{\vec{n}^{2}-\eta}-4\pi\Lambda\right].$ (5)
@@
Here, $\mu$ is the reduced mass of the two dineutrons, $B_{2n}$ is the dineutron binding energy, $\alpha=1/2$ for the dineutron, and $\Delta B_{2n}^{(L)}=B_{2n}^{(L)}-B_{2n}$ the finite volume correction. The last term is a topological factor that accounts for the finite-volume momentum-dependent effects Bour et al. (2011), where the summation over $\mathbf{k}$ includes all integer vectors.
@@ III Results and discussion
We employ a lattice spacing of $a=1.32$ fm, with $L$ ranging from 8 to 23, corresponding to physical box sizes from approximately 10 fm to 30 fm. For comparison, we adopted an interaction with SU(4) symmetry to fit the 1S0 scattering phase shifts. Details regarding its specific form and parameters are provided in the supplementary material.
@@
Figure 1: Ground-state energies of the $4n$ system calculated using NLEFT with both N3LO and SU(4) interactions, and the $2n$ system calculated using the Lanczos eigenvector method with the SU(4) interaction, as functions of the lattice size $L$. The error bars of NLEFT results indicate the uncertainty from the extrapolation in Eq. (S1).
@@
Fig. 1 shows the calculated energy of the tetraneutron system at different $L$ values, using both the full N3LO interaction ($a=1.32$ fm) and the simple SU(4) interaction ($a=1.64$ fm). We also indicate candidate resonance energies from experimental studies: $E=2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}\,$MeV from Ref. Duer and others (2022) and $E=0.83\pm 0.65\text{(stat.)}\pm 1.25\text{(sys.)}\,$MeV from Ref. Kisamori and others (2016). For comparison, the energy of the $2n$ system at different $L$ values using the SU(4) and N3LO interactions is computed via the Lanczos eigenvector method Lanczos (1950), which avoids Monte Carlo and extrapolation uncertainties.
@@
Both the N3LO and SU(4) interactions yield similar results for the $4n$ system, showing strongly repulsive behavior at small $L$ due to Pauli blocking, with energies gradually decreasing as $L$ increases. This consistency confirms earlier findings Lazauskas and Carbonell (2005); Hiyama et al. (2016); Gandolfi et al. (2017); Shirokov et al. (2016); Fossez et al. (2017); Deltuva (2018); Li et al. (2019); Marqués and Carbonell (2021) that the tetraneutron system is insensitive to details of the nuclear interaction as long as it is realistic. The smooth, continuous decrease of the $4n$ energy lacks the characteristic plateau formation expected for resonances in finite-volume methods Hazi and Taylor (1970); Zhang et al. (2008), suggesting a behavior typical of a non-bound system that does not form a resonance. Furthermore, we have examined the energy derivatives and find no evidence of resonant behavior in the tetraneutron system 1.
@@
Recent studies of the $4n$ energy distribution using the 8He($p,p^{4}$He)$4n$ reaction model suggest that the sharp low-energy peak observed by Duer and others (2022) can be explained by dineutron-dineutron correlations Lazauskas et al. (2023). In a finite volume, the $2n$ subsystem becomes quasi-bound in the sense that the lowest $S$-wave level is shifted to negative energy relative to the two-neutron threshold, even though the infinite-volume $nn$ system corresponds to a virtual state. This motivates us to study 2n2n-$2n$ scattering phase shifts using Lüscher’s finite-volume method, following Ref. Elhatisari et al. (2017). In this context, we employ a finite-volume dineutron approximation, which means that the two-neutron subsystem is treated as an effective weakly bound dimer inside the box, characterized by a large scattering length.
@@
Figure 2: Neutron-neutron ($n$-$n$) and neutron-dineutron (nn-$2n$) scattering phase shifts calculated using Lüscher’s finite-volume method with the SU(4) interaction at a lattice spacing a=1.64a=1.64\,fm. Points represent lattice data for different $B_{2n}$ values, and lines show fits based on the effective range expansion.
@@
Before turning to 2n2n-$2n$ scattering, we validate this finite-volume strategy by studying $n$-$n$ and nn-$2n$ scattering. Since at low energies and large separations the interaction details are then suppressed, we consider the zero range leading interaction here. We tune the coupling constant of the zero-range neutron-neutron interaction so that the two-neutron subsystem is close to threshold, and we parameterize this proximity by an effective “binding” scale $B_{2n}$ spanning $0.01$ to $0.09$ MeV. We then use the resulting couplings in the two-neutron and three-neutron systems and compute the finite-volume scattering energy levels, from which we extract the phase shifts using Lüscher’s method, as shown in Fig. 2. Here we define all quantities in lattice units, i.e., physical units multiplied by the corresponding power of $a$.
@@
The lattice data (points) and effective-range-expansion fits (lines) show that $p\cot\delta$ versus $p$ collapses onto the expected universal trend for all systems except at $B_{2n}=0.01$ MeV. This behavior is related to the fact that for short-range interactions with a large two-body scattering length, low-energy observables become insensitive to microscopic details of the interaction and are governed instead by universal large-scattering-length physics associated with the unitarity (conformal) fixed point Braaten and Hammer (2006). In this limit, dimensionless combinations such as $a\,p\cot\delta$ versus apa\,p approach universal functions, and ratios of scattering lengths (e.g., $a_{n\text{-}2n}/a_{n\text{-}n}$) tend toward constants, up to subleading range and regulator effects.
@@
Using the effective-range expansion, we extract the ratio of neutron-dineutron to neutron-neutron scattering lengths, $a_{n\text{-}2n}/a_{n\text{-}n}$. The results, summarized in Table 1, show a nearly constant ratio for most cases, in agreement with lattice calculations $1.176(6)$ Elhatisari et al. (2017) and semi-analytic continuum calculations $1.1791(2)$ Bedaque and van Kolck (1998); Bedaque et al. (1998); Gabbiani et al. (2000); Bedaque et al. (2003). We have verified that the results are robust against variations in lattice spacing from $a=1.32$ to $1.97$ fm 1.
Since Lüscher’s relation is derived for elastic scattering, another important consistency check is that the finite-volume levels used to extract nn-$2n$ phase shifts are not significantly affected by the opening of the breakup channel ($n+2n\to n+n+n$). In the finite-volume dimer picture, the relevant inelastic scale is set by the separation energy of the quasi-bound $2n$ level, which we characterize by $B_{2n}$. This corresponds to a characteristic momentum scale $\sqrt{mB_{2n}}$, with $m$ the nucleon mass. For our weakest binding this scale is small, nevertheless, even when the extracted momenta are larger than this estimate, we do not observe systematic departures from the effective-range-expansion description or anomalous volume dependence in $p\cot\delta$. This indicates that breakup (inelastic) effects are numerically negligible for the particular states and volumes analyzed here, and supports the use of the finite-volume dineutron approximation as an intermediate step toward $2n2n$ scattering.
Table 1: Ratio of neutron-dineutron to neutron-neutron scattering lengths for different dineutron binding energies $B_{2n}$ at a lattice spacing $a=1.64$ fm.
$B_{2n}$ $a_{n-2n}/a_{n-n}$ error
@@
We now turn to the extraction of the $S$-wave 2n2n-$2n$ phase shifts from our finite-volume spectra. For the realistic N3LO chiral EFT Hamiltonian of Ref. Elhatisari and others (2024), the infinite-volume two-neutron system is not bound, but it corresponds to a virtual state. Nevertheless, in the periodic boxes used here the lowest nnnn $S$-wave level is shifted below the two-neutron threshold and can be treated as a shallow finite-volume dimer (a quasi-bound dineutron). This provides a practical intermediate description in which the four-neutron finite-volume levels are interpreted as two composite dimers in relative motion, and Lüscher’s relation can be used to map the discrete energies to elastic 2n2n-$2n$ phase shifts.
@@
For each box size $L$ we compute the four-neutron energy $E_{4n}^{(L)}$ and the two-neutron energy $E_{2n}^{(L)}$ in the same periodic volume. The latter defines the effective finite-volume “binding” of the quasi-bound $2n$ level, $B_{2n}^{(L)}\equiv-E_{2n}^{(L)}$. To extract the relative momentum $p$ of the two-dineutron system we invert the finite-volume dispersion relation for composite dineutrons, Eq. (6). This inversion requires as input the infinite-volume dineutron binding energy $B_{2n}$. For the N3LO interaction this scale tends to zero as $L\!\to\!\infty$ and cannot be determined directly from our largest volumes. We therefore treat $B_{2n}$ as an auxiliary near-threshold parameter and scan representative values $B_{2n}=0.01$$0.09$ MeV (all below the smallest $B_{2n}^{(L)}$ encountered at our largest $L$), using the resulting spread as an estimate of the residual systematic uncertainty associated with employing the finite-volume dineutron picture when the underlying $nn$ pole is virtual.
@@
Figure 3: Dineutron-dineutron $S$-wave scattering phase shift $\delta$ as a function of relative momentum $p$ extracted from finite-volume $4n$ and $2n$ energies using Lüscher’s method and the composite-dimer finite-volume relation in Eq. (6), for the N3LO chiral EFT interaction of Ref. Elhatisari and others (2024). Different symbols correspond to the scanned near-threshold values of $B_{2n}$ (in MeV), and the shaded band indicates the momentum range associated with $L\simeq 20$$15$ fm.
@@
The resulting 2n2n-$2n$ $S$-wave phase shifts are shown in Fig. 3. At the smallest extracted relative momenta (largest boxes), the phase shifts show the largest sensitivity to the auxiliary near-threshold input $B_{2n}$ used in the inversion of Eq. (6). While some of the $B_{2n}$ values yield slightly negative $\delta(p)$ in this regime, the low-momentum trend systematically bends toward $\delta(p)\to 0$ as $B_{2n}\to 0$, as expected for short-range interactions. We therefore do not draw a firm conclusion from the sign of the lowest-momentum points in Fig. 3. As the momentum increases, $\delta(p)$ rises, becomes positive, and exhibits a shallow maximum of order $10^{\circ}$ around $p\simeq 60$ MeV, before decreasing again and turning negative at the highest momenta shown. The dependence on the scanned $B_{2n}$ values is mild over most of the range and is most pronounced only at the smallest momenta, where the dineutron size is largest and finite-volume effects are maximally amplified.
@@
In the intermediate window $p\simeq 55$$100$ MeV, $\delta(p)$ displays a clear non-monotonic structure, indicating a weak attractive feature. However, we do not observe the characteristic rapid rise through $90^{\circ}$ that would signal a narrow elastic resonance. The shaded band in Fig. 3 highlights $p\simeq 60$$84$ MeV, which corresponds to $L\simeq 20$$15$ fm and (for the same volumes) to confined $4n$ energies in the range $E_{4n}\simeq 1.7$$3.3$ MeV (cf. Fig. 1), overlapping the experimentally reported correlated $4n$ peak at $2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}$ MeV.

### v2

@@ Abstract
The nature of the tetraneutron ($4n$) system remains a pivotal question in nuclear physics. We investigate the $4n$ system using nuclear lattice effective field theory in finite volumes with a lattice size up to $L=30$ fm, employing both a high-precision N3LO interaction and a simplified SU(4) symmetric one. The ground-state energy is found to decrease smoothly with increasing box size, showing no plateau characteristic of a resonance. We further compute the dineutron-dineutron scattering phase shift using Lüscher’s finite-volume method. At the smallest relative momenta, the extracted $2n2n$ $S$-wave phase shift is small, consistent with a weak interaction in the dilute limit. At intermediate momenta, it exhibits a weak attraction with a peak of approximately $10^{\circ}$ at relative momentum of 60–84 MeV. While this structure does not constitute a resonance, the corresponding confined $4n$ energy of 1.7–3.3 MeV lies close to the experimentally observed low-energy peak.
@@ I Introduction
Bound nuclear systems conventionally form when protons and neutrons coalesce through the strong interaction. However, the combination of attractive strong interactions and the absence of Coulomb repulsion makes the existence of pure neutron nuclei a plausible concept, a quest that dates back to the early 1960s ogloblin1989nuclei. Confirming or ruling out the existence of such neutral nuclei, whether as bound or resonant states, is crucial for advancing our understanding of nuclear physics, the strong interaction, universal fermion properties in the unitary limit Elhatisari:2016hui; Braaten:2004rn, and astrophysical objects Matsuki:2024ios; Ivanytskyi:2019ynz.
@@
While a bound $4n$ nucleus is generally believed to be excluded Lazauskas:2022mvq; Marques:2021mqf; Bertulani:2002px, the interpretation of these signals as a resonance remains highly debated. A variety of theoretical approaches, including Jost function analysis in the complex momentum plane Sofianos:1997nn, the hyperspherical harmonic method Grigorenko:2004, Faddeev-Yakubovsky equations Lazauskas:2005ig; Hiyama:2016nwn; Carbonell:2017dfl, the Gaussian expansion method Hiyama:2016nwn; Carbonell:2017dfl, the no-core Gamow shell model and density matrix renormalization group Fossez:2016dch, the $4n$ response function Lazauskas:2017cfi, Alt, Grassberger, and Sandhas (AGS) equations Deltuva:2018xoa; Deltuva:2019mnv, and adiabatic hyperspherical methods Higgins:2020avy; Higgins:2020pbe converge on the absence of a $4n$ resonance. The low-energy peak observed by Duer:2022ehf has been explained as a consequence of dineutron-dineutron correlations through a reaction model study Lazauskas:2022mvq. Reactions with multineutron final states using a nonrelativistic conformal field theory are investigated in Hammer:2021zxb. Recent tensor-optimized antisymmetrized molecular dynamics studies suggest that a $4n$ resonance at the nuclear surface may be influenced by core attraction Wan:2025itw. For further insight into multineutron correlations in light nuclei, see Zhang:2025uin.
@@
In contrast, several many-body calculations have reported a low-lying $4n$ resonance with energies and widths close to the experimental values. Some of these results are obtained by introducing an external confining potential and extrapolating to the zero-trap limit, e.g., in quantum Monte Carlo Pieper:2003dc; Gandolfi:2016bth and no-core Gamow shell–model studies Li:2019pmg. Other approaches do not employ an explicit trap, such as the single-state harmonic oscillator representation of scattering equations extension of the no-core shell model Shirokov:2016ywq. However, as discussed in Refs. Deltuva:2019ngx; Ishikawa:2020bcs, threshold effects are critical and must be carefully accounted for in such extrapolations. For a comprehensive review of previous efforts, see Marques:2021mqf.
@@
To further elucidate the nature of the $4n$ signal reported in Duer:2022ehf, additional theoretical investigations are needed. In this work, we study the $4n$ system using nuclear lattice effective field theory (NLEFT) Lee:2025req; Lahde:2019npb, a method that has proven very successful in providing a unified description of nuclear phenomena on a discretized lattice, including scattering Elhatisari:2025fyu; Wu:2025fkn; Elhatisari:2016hui; Elhatisari:2015iga, clustering Epelbaum:2011md; Epelbaum:2012qn; Shen:2022bak; Shen:2024qzi, nuclear matter Tong:2025fzv; Tong:2024jvs; Ma:2023ahg; Ren:2023ued; Lu:2019nbg, and the structure of light and heavy nuclei Hildenbrand:2025voq; Ren:2025vpe; Song:2025ofd; Niu:2025uxk; Wang:2025swg; Zhang:2024wfd; Hildenbrand:2024ypw; Meissner:2023cvo; Sarkar:2023qjn; Lu:2021tab; Shen:2021kqr. We calculate the ground-state energy of $4n$ in finite volumes up to a lattice size of $L\simeq 30$ fm. No external trap is applied, as the finite volume itself provides the necessary confinement for the dilute system.
@@ II Formalism
with $\mathbf{n}=(n_{x},n_{y},n_{z})$ is the lattice coordinate vector, with each component ranging from 0 to L1L-1, and $\mathbf{k}$ is the lattice momentum vector. The initial state $|\Psi_{0}\rangle$ is constructed as a Slater determinant from the two lowest momentum states, each with spin up and down. Furthermore, we perform an angular momentum projection onto the $A_{1}^{+}$ irreducible representation of the cubic group to project out the $0^{+}$ state and accelerate the Euclidean time evolution Johnson:1982yq; Lu:2014xfa. The ground state energy at a given projection time $\tau$ is calculated as
@@
It is well known that the tetraneutron system lacks bound subsystems. However, when confined in a finite volume, the dineutron becomes quasibound. Following the method for dimer-dimer scattering on a lattice Elhatisari:2016hui, we compute the $2n2n$ scattering phase shifts using Lüscher’s formula. It relates the two-body energy levels in a periodic cubic box to the elastic scattering phase shifts Luscher:1986pf; Luscher:1990ux:
@@
with $S(\eta)$ the regulated three-dimensional $\zeta$ function
$S(\eta)=\lim_{\Lambda\to\infty}\left[\sum_{\vec{\bm{n}},\,|\vec{\bm{n}}|\leq\Lambda}\frac{1}{\vec{\bm{n}}^{2}-\eta}-4\pi\Lambda\right].$ (5)
@@
Here, $\mu$ is the reduced mass of the two dineutrons, $B_{2n}$ is the dineutron binding energy, $\alpha=1/2$ for the dineutron, and $\Delta B_{2n}^{(L)}=B_{2n}^{(L)}-B_{2n}$ the finite volume correction. The last term is a topological factor that accounts for the finite-volume momentum-dependent effects Bour:2011ef, where the summation over $\mathbf{k}$ includes all integer vectors.
@@ III Results and discussion
We employ a lattice spacing of $a=1.32$ fm, with $L$ ranging from 8 to 23, corresponding to physical box sizes from approximately 10 fm to 30 fm. For comparison, we adopted an interaction with SU(4) symmetry to fit the 1S0 scattering phase shifts. Details regarding its specific form and parameters are provided in Supplementary Material.
@@
Figure 1: Ground-state energies of the $4n$ system calculated using NLEFT with both N3LO and SU(4) interactions, and the $2n$ system calculated using the Lanczos eigenvector method with the SU(4) interaction, as functions of the lattice size $L$. The error bars of NLEFT results indicate the uncertainty from the extrapolation SM.
@@
Fig. 1 shows the calculated energy of the tetraneutron system at different $L$ values, using both the full N3LO interaction ($a=1.32$ fm) and the simple SU(4) interaction ($a=1.64$ fm). We also indicate candidate resonance energies from experimental studies: $E=2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}\,$MeV from Ref. Duer:2022ehf and $E=0.83\pm 0.65\text{(stat.)}\pm 1.25\text{(sys.)}\,$MeV from Ref. Kisamori:2016jie. For comparison, the energies of the $2n$ system at different $L$ values using the SU(4) and N3LO interactions are computed via the Lanczos eigenvector method Lanczos:1950zz, which eliminates Monte Carlo and extrapolation uncertainties.
@@
Both the N3LO and SU(4) interactions yield similar results for the $4n$ system, showing strongly repulsive behavior at small $L$ due to Pauli blocking, with energies gradually decreasing as $L$ increases. This consistency confirms earlier findings Lazauskas:2005ig; Hiyama:2016nwn; Gandolfi:2016bth; Shirokov:2016ywq; Fossez:2016dch; Deltuva:2018xoa; Li:2019pmg; Marques:2021mqf that the tetraneutron system is insensitive to details of the nuclear interaction as long as it is realistic. The smooth, continuous decrease of the $4n$ energy lacks the characteristic plateau formation expected for resonances in finite-volume methods Hazi:1970; Zhang:2007hn, suggesting a behavior typical of a non-bound system that does not form a resonance. Furthermore, we have examined the energy derivatives and find no evidence of resonant behavior in the tetraneutron system SM. Calculations at large box sizes are computationally demanding and yield larger Monte Carlo uncertainties. Nevertheless, we extend our N3LO calculations up to $L\simeq 30$ fm, reaching energies down to $E\simeq 0.83$ MeV, i.e. within the range of experimental candidate tetraneutron resonance energies. In the L=25L=253030 fm window the energy decreases monotonically and does not become volume independent within uncertainties, so no plateau is observed.
@@
Recent studies of the $4n$ energy distribution using the 8He($p,p^{4}$He)$4n$ reaction model suggest that the sharp low-energy peak observed by Duer:2022ehf can be explained by dineutron-dineutron correlations Lazauskas:2022mvq. In a finite volume, the $2n$ subsystem becomes quasibound in the sense that the lowest $S$-wave level is shifted to negative energy relative to the two-neutron threshold, even though the infinite-volume nnnn system corresponds to a virtual state. This motivates us to study $2n2n$ scattering phase shifts using Lüscher’s finite-volume method, following Ref. Elhatisari:2016hui. In this context, we employ a finite-volume dineutron approximation, which means that the two-neutron subsystem is treated as an effective weakly bound dimer inside the box, characterized by a large scattering length.
@@
Figure 2: Neutron–neutron (nnnn) and neutron–dineutron (nn2n2n) scattering phase shifts calculated using Lüscher’s finite-volume method with the SU(4) interaction at a lattice spacing a=1.64a=1.64\,fm. Points represent lattice data for different $B_{2n}$ values, and lines show fits based on the effective range expansion.
@@
Before turning to $2n2n$ scattering, we validate this finite-volume strategy by studying nnnn and nn2n2n scattering. Since at low energies and large separations the interaction details are then suppressed, we consider the zero-range leading interaction here. We tune the coupling constant of the zero-range neutron-neutron interaction so that the two-neutron subsystem is close to the threshold, and we parameterize this proximity by an effective “binding” scale $B_{2n}$ spanning $0.01$ to $0.09$ MeV. We choose this range to explore shallow finite-volume dimers close to the nnnn threshold. The lower end $B_{2n}=0.01$ MeV deliberately probes the extreme near-threshold regime, where finite-volume and lattice-spacing effects are amplified (see discussion below), while the upper end $B_{2n}=0.09$ MeV remains sufficiently close to threshold to stay in the large-scattering-length regime but provides a practical range for systematic checks.
We then use the resulting couplings in the two-neutron and three-neutron systems and compute the finite-volume scattering energy levels, from which we extract the phase shifts using Lüscher’s method, as shown in Fig. 2. Here we define all quantities in lattice units, i.e., physical units multiplied by the corresponding power of $a$.
@@
The lattice data (points) and effective-range-expansion fits (lines) show that $p\cot\delta$ versus $p$ collapses onto the expected universal trend for all systems except at $B_{2n}=0.01$ MeV. This behavior is related to the fact that for short-range interactions with a large two-body scattering length, low-energy observables become insensitive to microscopic details of the interaction and are governed instead by universal large-scattering-length physics associated with the unitarity (conformal) fixed point Braaten:2004rn. In this limit, dimensionless combinations such as $a\,p\cot\delta$ versus apa\,p approach universal functions, and ratios of scattering lengths (e.g., $a_{n\text{-}2n}/a_{n\text{-}n}$) tend toward constants, up to subleading range and regulator effects.
@@
Using the effective-range expansion, we extract the ratio of neutron-dineutron to neutron-neutron scattering lengths, $a_{n\text{-}2n}/a_{n\text{-}n}$. The results, summarized in Table 1, show a nearly constant ratio for most cases, in agreement with lattice calculations $1.176(6)$ Elhatisari:2016hui and semianalytic continuum calculations $1.1791(2)$ Bedaque:1997qi; Bedaque:1998mb; Gabbiani:1999yv; Bedaque:2002yg. We have verified that the results are robust against variations in lattice spacing from $a=1.32$ to $1.97$ fm SM.
Since Lüscher’s relation is derived for elastic scattering, another important consistency check is that the finite-volume levels used to extract nn2n2n phase shifts are not significantly affected by the opening of the breakup channel ($n+2n\to n+n+n$). In the finite-volume dimer picture, the relevant inelastic scale is set by the separation energy of the quasibound $2n$ level, which we characterize by $B_{2n}$. This corresponds to a characteristic momentum scale $\sqrt{mB_{2n}}$, with $m$ the nucleon mass. For our weakest binding this scale is small; nevertheless, even when the extracted momenta are larger than this estimate, we do not observe systematic departures from the effective-range-expansion description or anomalous volume dependence in $p\cot\delta$. This indicates that breakup (inelastic) effects are numerically negligible for the particular states and volumes analyzed here, and supports the use of the finite-volume dineutron approximation as an intermediate step toward $2n2n$ scattering.
Table 1: Ratio of neutron–dineutron to neutron–neutron scattering lengths for different dineutron binding energies $B_{2n}$ at a lattice spacing $a=1.64$ fm. The column of uncertainty indicates the statistical uncertainty from the fits to the effective range expansion.
$B_{2n}$ $a_{n-2n}/a_{n-n}$ Uncertainty
@@
We now turn to the extraction of the $S$-wave $2n2n$ phase shifts from our finite-volume spectra. For the realistic N3LO chiral EFT Hamiltonian of Ref. Elhatisari:2022zrb, the infinite-volume two-neutron system is not bound, but it corresponds to a virtual state. Nevertheless, in the periodic boxes used here the lowest nnnn $S$-wave level is shifted below the two-neutron threshold and can be treated as a shallow finite-volume dimer (a quasibound dineutron). This provides a practical intermediate description in which the four-neutron finite-volume levels are interpreted as two composite dimers in relative motion, and Lüscher’s relation can be used to map the discrete energies to elastic $2n2n$ phase shifts.
@@
For each box size $L$ we compute the four-neutron energy $E_{4n}^{(L)}$ and the two-neutron energy $E_{2n}^{(L)}$ in the same periodic volume. The latter defines the effective finite-volume “binding” of the quasibound $2n$ level, $B_{2n}^{(L)}\equiv-E_{2n}^{(L)}$. To extract the relative momentum $p$ of the two-dineutron system we invert the finite-volume dispersion relation for composite dineutrons [Eq. (6)]. This inversion requires as input the infinite-volume dineutron binding energy $B_{2n}$. For the N3LO interaction this scale tends to zero as $L\!\to\!\infty$ and cannot be determined directly from our largest volumes. We therefore treat $B_{2n}$ as an auxiliary near-threshold parameter and scan representative values $B_{2n}=0.01$$0.09$ MeV (all below the smallest $B_{2n}^{(L)}$ encountered at our largest $L$), using the resulting spread as an estimate of the residual systematic uncertainty associated with employing the finite-volume dineutron picture when the underlying nnnn pole is virtual.
@@
Figure 3: Dineutron–dineutron $S$-wave scattering phase shift $\delta$ as a function of relative momentum $p$ extracted from finite-volume $4n$ and $2n$ energies using Lüscher’s method and the composite-dimer finite-volume relation in Eq. (6), for the N3LO chiral EFT interaction of Ref. Elhatisari:2022zrb. Different symbols correspond to the scanned near-threshold values of $B_{2n}$ (in MeV), and the shaded band indicates the momentum range associated with $L\simeq 20$$15$ fm.
@@
The resulting $2n2n$ $S$-wave phase shifts are shown in Fig. 3. At the smallest extracted relative momenta (largest boxes), the phase shifts show the largest sensitivity to the auxiliary near-threshold input $B_{2n}$ used in the inversion of Eq. (6). While some of the $B_{2n}$ values yield slightly negative $\delta(p)$ in this regime, the low-momentum trend systematically bends toward $\delta(p)\to 0$ as $B_{2n}\to 0$, as expected for short-range interactions. We therefore do not draw a firm conclusion from the sign of the lowest-momentum points in Fig. 3. As the momentum increases, $\delta(p)$ rises, becomes positive, and exhibits a shallow maximum of order $10^{\circ}$ around $p\simeq 60$ MeV, before decreasing again and turning negative at the highest momenta shown. The dependence on the scanned $B_{2n}$ values is mild over most of the range and is most pronounced only at the smallest momenta, where the dineutron size is largest and finite-volume effects are maximally amplified.
@@
In the intermediate window $p\simeq 55$$100$ MeV, $\delta(p)$ displays a clear nonmonotonic structure, indicating a weak attractive feature. However, we do not observe the characteristic rapid rise through $90^{\circ}$ that would signal a narrow elastic resonance. The shaded band in Fig. 3 highlights $p\simeq 60$$84$ MeV, which corresponds to $L\simeq 20$$15$ fm and (for the same volumes) to confined $4n$ energies in the range $E_{4n}\simeq 1.7$$3.3$ MeV (cf. Fig. 1), overlapping the experimentally reported correlated $4n$ peak at $2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}$ MeV.
Our results are qualitatively consistent with previous investigations of $2n2n$ scattering phase shifts using rigorous solutions of FY and AGS equations Deltuva:2019mnv, which found that at physical interaction strengths, a bound tetraneutron state evolves into a virtual state rather than a resonance. While we identify a weak nonmonotonic attraction at intermediate momenta in the current framework, the phase shift remains well below $90^{\circ}$, supporting the conclusion that no observable narrow resonance exists at the physical point.